OpenStudy (anonymous):

Can someone just help me with the first step of solving this please? 11-7p<-10or 8p+2<-46

OpenStudy (anonymous):

I think for the first part I add seven.

OpenStudy (anonymous):

no 11-7p<-10 subtract 11 -7p<-10-11 then flipe the sign and div by 7 p>-21/7 p>-3

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Minus 8?

OpenStudy (anonymous):

8p+2<-46 8p<-48 subtract 2 p<-48/8 divide 8

OpenStudy (anonymous):

or 2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

p<-6?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

when something has a coefficient you divide not subtract like / no coefficient. ax+c=d ^ coefficien therefore first subtract c ax=d-c then divide by coefficien x=(d-c)/a